The average revenue function for a commodity is p = 50 – 4q. Find edp when:

demand = 5 units ; price = Rs. 6. Find consumer’s surplus at price = Rs. 6



1
Expert's answer
2021-05-09T18:15:08-0400


p is the price and q quantituy

Average revenue = "\frac{Total revenue}{q}"  

Total revenue ="p\times q"  

Then 

Average revenue "=\frac{ Pq}{q} = p"

Therefore

Average revenue function is the demand function. 

We can write inverse demand function as: 

"p = 50 - 4q"

"q =\frac {50-p}{4}"

Price elasticity of demand "=\frac{ dq}{dp} ×\frac{ p}{q}"

"\frac{dq}{dp}= \frac{-1}{4}"

Price elasticity of demand at "q = 5" is equal to: 

"P = 50 - 4(5) = 30."

"\frac{dq}{dp}= \frac{-1}{4}."

Price elasticity of demand "= \frac{-1}{4} ×\frac{ 30}{5} = -1.5"

At price "= 6"  

"q =\frac{50-6}{4} =\frac{ 44}{4} = 11"

Price elasticity of demand"=\frac{ -1}{4} ×\frac{6}{11}

= -0.136"


Price "= 6"

Then"q = 11"

P when q q "= 0 is 50 - 4(0) = 50"

Consumer surplus"= \frac{1}{2} × (50-6) × 11"

Consumer surplus "=\frac{1}{2} × 44 × 11"

Consumer surplus "= 242"